Operation Summary for July 16, 2010 - July 19, 2010 |
The Integrated Luminosity for the period from 7/12/10 to 7/19/10 was 44.63 pb-1. Friday July 16 The day shift began with Operations monitoring the Tevatron (TeV) store 7979 (157.58E30) and the Recycler stash (96.4E10), with the Antiproton Source (Pbar) stacking (17.6E10), with the Linac, Booster and the Main Injector (MI) operating normally, with the MiniBooNE and Meson FTBF T-977 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, various machine personnel conducted machine studies. At 3:33 PM, the stack disappeared; operators investigated and discovered a power supply (A:SQ100) problem in the Accumulator had caused the loss. During the evening shift, Operations terminated store 7979 at 9:30 PM and put the TeV into HEP shot setup. TeV personnel conducted studies during shot setup. A Linac RF station (LRF1) developed an RF problem, which Linac personnel fixed, but this delayed shot setup. Operations disabled the T-977 beam at midnight. Saturday July 17 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (87.2E10), with Pbar stacking (51.3E10), and with MiniBooNE taking beam. During the midnight shift, Operations disabled stacking at 12:27 AM. Operations established store 7980 at 12:30 AM with an average initial luminosity of 229.3E30. Operators turned off all power supplies in MI, Switchyard (SY), Meson, MiniBooNE, and Booster, with the exception of the Booster Gradient Magnet Power Supply. Recycler Personnel conducted studies. LRF1 tripped off; operators left the station off per instructions. During the day shift, Pbar personnel conducted studies until 11:25 AM – operators then began preparing the Pbar enclosures and equipment for shutdown. Recycler studies ended at noon – operators began preparing the Recycler and the Pelletron for shutdown. Interlock technicians conducted P1 transfer line safety system tests from 8:57 AM to 11:59 AM. During the evening shift, operators prepared Linac and Booster for shutdown. Operators prepared the SY and Meson enclosures for shutdown. Sunday July 18 The midnight shift began with Operations monitoring store 7980 and the stash, with Pbar stacking, and with and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7980 at 12:06 AM. The Cryo System Coordinator reported at 1:27 AM that CHL personnel had started the TeV sector A3 cryo system warm up. Operators began preparing the TeV for shutdown. During the day shift, Safety techs recored the door and gate locks in Booster, MI, and the TeV. Operations conducted training. During the evening shift, some shutdown activities began. Monday July 19 The midnight shift began with the accelerator complex prepared for shutdown. During the midnight shift, Operations reported two vacuum bursts in the MI-40 region; operators brought the ion pumps back online. More shutdown activities started. The Plans for Monday The plans for today are to begin a safe and efficient repair period. For most areas, this shutdown will last for four weeks. |
|
More Information Comments and Suggestions |
| last modified 07/19/2010 email Fermilab |
| Security, Privacy, Legal | |